Influence of manufacturing process on Cavitation Erosion on CoCrWMoCFeNiSiMn (Stellite 1) alloys

Vishakh Pradeep Kumar 2025-08-14 Thu

Agenda

• Introduction

• Aims

• Methodology

• Results & Discussion • Conclusion

Introduction

Cavitation Erosion

• What is cavitation? Collapse of bubbles and the resulting high-frequency high-pressure shock waves. Caused by fluid pressure dropping to vapor pressure, which is particularly common with high fluid flow speeds [1].

• Why does it matter? Cavitation erosion leads to removal of material, crack growth, and part failure. Affects turbine blades, pump impellers, valves, stirrers, etc.

Example (A screenshot)

Stuff

Stellite mind map

This slide consists of some text with a number of bullet points:

Why

4

A simple slide

This slide consists of some text with a number of bullet points:

• the first, very @important@, point!

• the previous point shows the use of the special markup which translates to the

Beamer specific alert command for highlighting text. The above list could be numbered or any other type of list and may include sub-lists.

A more complex slide

This slide illustrates the use of Beamer blocks. The following text, with its own headline, is displayed in a block: Theorem (Org mode increases productivity)

• org mode means not having to remember L^{A}TEX commands.

• it is based on ascii text which is inherently portable.

• Emacs!

Two columns

• this slide consists of two columns

• the first (left) column has no heading and consists of text

• the second (right) column has an image and is enclosed in an example block Example (A screenshot)

Stuff

Aims Methodology